OpenStudy (anonymous):

solve the following system of equations. y=x+4-2x=10 my answers include (5,1), (-5,-1) (5,1) (4,4)

OpenStudy (anonymous):

What you have to do first is put the equation into slope intercept form: y=mx+b y = x + 4 - 2x = 10 Could that possibly be + 10?

OpenStudy (anonymous):

no that is not one of my answers. I am confused and tired.

OpenStudy (anonymous):

I wasn't giving that as an answer. I was making sure that your notation was correct.

OpenStudy (anonymous):

Yes that right but I need the aswer

OpenStudy (anonymous):

I will give a medal!!!

OpenStudy (anonymous):

So it's y = x + 4 - 2x + 10 Simplify the x and -2x Simplify the 4 and 10 y = -1x + 14 Are you with me so far?

OpenStudy (anonymous):

Sortof

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

I only simplified the terms. Now to find the value of "y," I have to make x = 0 y = 1 Now in order to find the value of "x," I only have to solve x = 5 Your answer is (5,1) Does that make sense?